OpenStudy (alante):

What is the solution of the system? y = 3x - 6 x + 2y = 2 A. (-3, 0) B. (2, 0) C. (0, 2) D. (0, -3)

OpenStudy (jjcoolja12):

Problem: Solve y=3x−6;x+2y=2 Steps: I will try to solve your system of equations. y=3x−6;x+2y=2 Step: Solvey=3x−6for y: Step: Substitute3x−6foryinx+2y=2: x+2y=2 x+2(3x−6)=2 7x−12=2(Simplify both sides of the equation) 7x−12+12=2+12(Add 12 to both sides) 7x=14 7x 7 = 14 7 (Divide both sides by 7) x=2 Step: Substitute2forxiny=3x−6: y=3x−6 y=(3)(2)−6 y=0(Simplify both sides of the equation) Answer: x=2 and y=0

OpenStudy (jjcoolja12):

so it's b